OpenStudy (anonymous):

PLEASE HELP I AM STUCK ON THIS PROBLEM !!!!!!!!!!!!!!!!! The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 3 miles west and 2 miles south of the City Center. The park is 4 miles east and 5 miles north of the Center. How far is it from the mall to the park to the nearest tenth of a mile?

OpenStudy (wolfe8):

Sketching what info you have will help a lot. You'll usually end up with just a geometry problem. Let me know if you need help with that.

OpenStudy (anonymous):

Find the distance from (-5,-3) to (4,4) Distance= Sqrt[(-5-4)^2+(-3-4)^2]=Sqrt(81+49)= Sqrt(130). Without a calculator handy, the square root of 130 will be more than 11 (because 11*11=121) and less than 12 (12*12=144) but a little closer to 11. D. 11.4 miles (on a calculator 11.40175425099138)

OpenStudy (anonymous):

thkx

OpenStudy (anonymous):

your welcome

OpenStudy (jdoe0001):

see where the distances, where you can use the pythagorean theorem to get the "blue line" or hypotenuse